OpenStudy (migitmack):

Rachel moved into a new apartment m months ago and paid a $1415 deposit. She pays a monthly rent of $750. How many months has she lived in her apartment if she has paid $8,165? Is it about 10 months?

OpenStudy (migitmack):

My answer was a bit over 8,165$

OpenStudy (migitmack):

i divided 8,165 with 750 and got 10.86

OpenStudy (ragingsquirrel):

What about the deposit?

OpenStudy (anonymous):

K well there's 10 months and she pays $765 each month. 765*10=7650. Now we just add that with the $1415 deposit and we get $8915. I would consider it more to the 11 month side than the 10 month.

OpenStudy (anonymous):

It's acutally 11 months^ 8165+750=8915

OpenStudy (migitmack):

yeah

OpenStudy (migitmack):

wait

OpenStudy (migitmack):

thats wrong I got 9k

OpenStudy (migitmack):

OK I GOT IT NOW THX

OpenStudy (anonymous):

K look. To check to see if it's 10 months, we do 765*10=7650. That's her monthly payments. Now we add it with the deposit, and that equals $8915, which means that she's been there more than 10 months.

OpenStudy (migitmack):

the answer is about 11 months

OpenStudy (anonymous):

Yep :D

OpenStudy (anonymous):

First take the deposit off the amount paid which would equal 6750. Then divide 6750 by 750 and you would get 9 months